OpenStudy (anonymous):

Part 1: Solve each of the quadratic equations below. Show your work. (3 points) x2 − 25 = 0 and x2 = 2x + 15 Part 2: Describe what the solution(s) represent to the graph of each. (2 points) Part 3: How are the graphs alike? How are they different? (2 points)

OpenStudy (anonymous):

first one \[x^2-25=0\] \[x^2=25\] \[x=\pm\sqrt{25}=\pm5\]

OpenStudy (anonymous):

which really means \(x=-5\) or \(x=5\)

OpenStudy (anonymous):

second one \[ x^2 = 2x + 15\] \[x^2-2x-15=0\] then factor

OpenStudy (anonymous):

Part 2: Describe what the solution(s) represent to the graph of each. the zeros are where the graph of \(y=x^2-25\) cross the \(x\) axis same for \(y=x^2-2x-15\)

OpenStudy (anonymous):

thx .[[[:

OpenStudy (anonymous):

Part 3: How are the graphs alike? How are they different?